Circular No. 6622

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 G1 Joe Montani, Lunar and Planetary Laboratory, reports the discovery

of a comet with the Spacewatch 0.9-m telescope at Kitt Peak. J. Larsen assisted with the observations on Apr. 9. The measurements below are by Montani (Apr. 9, showing a tail extending for 2′ from the coma in p.a. 285^{o}) and by J. V. Scotti (Apr. 10, coma diameter 10′′, tail extending 1′.3 in p.a. 284^{o}). On Apr. 10.122 UT, m_{2} = 21.4.

1997 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 9.13872  7^{h}16^{m}09.^{s}59  +19^{o}59′45′′.7  19.1
9.14671  7 16 09.91  +19 59 44.9  19.1
9.16504  7 16 10.53  +19 59 44.0  19.0
10.12171  7 16 44.12  +19 58 42.3
10.13250  7 16 44.49  +19 58 41.4  19.9
10.13897  7 16 44.76  +19 58 41.7  19.4

SUPERNOVA 1997bp IN NGC 4680 L. Wang and J. C. Wheeler, University of Texas, communicate: “Spec-

tropolarimetry of SN 1997bp was obtained on Apr. 7.2 UT with the 2.7-m telescope (+ SPOL) at McDonald Observatory. Preliminary reduction of the data (total exposure 4 hr) indicates polarization variation at a level of > 0.5 percent across the strong Si II 635.5-nm feature. Such a variation is indicative of intrinsic polarization of the supernova. More observations, particularly polarimetry, are strongly incouraged. For comparison, our data

on SN 1997bp can be accessed at http://tycho.as.utexas.edu/~lifan,
or by e-mail to lifan@astro.as.utexas.edu.”

SUPERNOVA 1997bq IN NGC 3147 W.-d. Li, Beijing Astronomical Observatory (BAO), reports that an

observation of NGC 3147 on Mar. 27 shows no object at the location of SN 1997bq (cf. IAUC 6616) to a limiting mag of 18.5. On Apr. 3, and unfiltered CCD frame shows the supernova at mag ∼ 18.0. On Apr. 9, photometric observations with the BAO 0.6-m reflector yield the following magnitudes for SN 1997bq (calibration stars from M67): B = 15.64 ± 0.04, V = 15.49 ± 0.03, R = 14.84 ± 0.03.

1997 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green